Show that the series , ∞Σ n=1 (x/1+n^2x^2) is uniformly convergent in [ k,1] for any k>0.

1
Expert's answer
2022-01-20T08:17:06-0500

If "x\in[k,1]" , "0<k<1" , then "\frac{x}{1+n^2x^2}\leq \frac{x}{n^2x^2}=\frac{1}{n^2x}\leq \frac{1}{n^2 k}".

The expression on the right hand does not depend on x and the series "\sum\limits_{n=1}^{\infty}\frac{1}{n^2 k}" is convergent, therefore the initial series is uniformly convergent on the segment [k,1].


In order to prove that the series is uniformly convergent for any "k>0" , it is sufficient to show that for any "N\in\mathbb{N}" there exists "x\in[0,1]" , "q> p> N" such that "\sum\limits_{n=p}^{q}\frac{x}{1+n^2x^2}\geq 1/5" .

Put x=1/N, p=N+1, q=2N. Then "nx\leq 2" for all "n=p, p+1, \dots, q" and

"\sum\limits_{n=p}^{q}\frac{x}{1+n^2x^2}\geq \sum\limits_{n=p}^{q}\frac{x}{1+2^2}=\sum\limits_{n=N+1}^{2N}\frac{1/N}{5}=1/5"

Therefore, the series "\sum\limits_{n=1}^{\infty}\frac{x}{1+n^2x^2}" is uniformly convergent for any "k>0" .